# Constrained Measurement Incompatibility from Generalised Contextuality of Steered Preparation

### v2

@@ VIII Summary and Conclusions
We note here that a number of studies have been made toward establishing a similar connection between incompatibility and various forms of quantum correlations, viz. nonlocality [13, 14] steering [22, 51], preparation and measurement noncontextuality [18, 63]. In this work, we have moved a step forward by establishing the connection between measurement incompatibility and generalised noncontextuality in any no-signalling operational theory, and thereby quantifying the maximum allowed incompatibility of a theory determined by the generalised contextuality.
@@
Furthermore, we have discussed the geometrical implications of the $N$-wise compatibility in a GPT. It has recently been proved in [20] that a GPT is simplex-embeddable if and only if the ontological model corresponding to the operational theory this GPT concerns is generalised noncontextual. We demonstrated that any set of Alice’s observables that is used to remotely prepare Bob’s states is compatible if and only if Bob’s local statistics can be described in terms of a simplex-embeddable GPT. In [66, 65], it is shown that any preparation whose corresponding statistics cannot be described in terms of a simplex-embeddable GPT can be made simplex-embeddable by allowing dephasing noise to the states. In contrast to Refs.[66, 65], where prepare-measure scenario was considered, we argued that in our bipartite Bell scenario, Alice’s unsharp measurements remotely prepares noisy states so that a simplex-embeddable GPT explanation exists (which is otherwise not simplex-embeddable). Obviously, the degree of unsharpness parameter needs to be bounded. In quantum theory, simplex-embeddability exists for $\eta\leq 1/\sqrt{N}$.
@@
In recent work [64, 45], it is argued that measurement incompatibility is not a prerequisite to reveal generalised contextuality in the preparation and measurement scenario. Although it apparently seems, there is no contradiction with the results of our paper. We claimed that the compatibility of Alice’s measurements used to steer Bob’s states is necessary and sufficient to reveal the generalised contextuality of the steered preparations. On the other hand, the claim of [64, 45] refers only to the connection between the measurements used to obtain the statistics from a single system and the possibility of revealing the generalised contextuality. Therefore, the domain of applicability of the results obtained in [64, 45] and this paper is different.
@@
Finally, we may consider a more general definition of the incompatibility of a theory in terms of the compatibility structure that exists in such a theory. In fact, it would be interesting to find connections between non-classical correlations exhibited by an arbitrary operational theory and other layers of non-classicality [67, 68] exist for measurements apart from the $N$-wise incompatibility. This would enable one to explore the uniqueness of quantum theory compared to other no-signaling theories. We note that, while discarding other class of incompatible theories in terms of our definition, we have assumed the superiority of generalized contextuality of quantum theory. It would be more intriguing if one can find a physical principle that can discard such theories and super-select the quantum theory. Such an attempt [69] was made which is particularly designed for the global joint measurability of sharp measurements through the Speaker’s principle. Studies along those direction could be an exciting avenue for future research and thus calls for further study.
@@
Acknowledgements:- S.M. acknowledges David Schmid for his comment on an earlier version of the manuscript. A.K.P acknowledges the support from the research grant SERB/MTR/2021/000908, Government of India.
@@ Appendix A Derivation of necessary condition for $N$-wise compatibility
$\ \ \ \ \ \ \ \ \ \ \ =\frac{1}{2^{N-1}}\langle\ (A_{1}+A_{2}+.....+A_{N})\ \otimes B_{1}\ \rangle,$ (47)
Further, by using Eq. (45)-(A), we can write,
$p(a_{1}=a_{2}=.....=a_{N}|A_{1},A_{2},....,A_{N})\geq\frac{1}{2^{N-1}}\lvert\langle\ (A_{1}+A_{2}+.....+A_{N})\ \otimes B_{1}\ \rangle\rvert$ (48)
@@
$p(a_{1}=-a_{2}=.....=a_{N}|A_{1},A_{2},....,A_{N})$$\geq$ $\frac{1}{2^{N-1}}\lvert\langle\ (A_{1}-A_{2}+A_{3}.....+A_{N})\ \otimes B_{2}\ \rangle\rvert$
$p(a_{1}=a_{2}=-a_{3}=.....=a_{N}|A_{1},A_{2},A_{3}....,A_{N})$$\geq$ $\frac{1}{2^{N-1}}\lvert\langle\ (A_{1}+A_{2}-A_{3}.....+A_{N})\ \otimes B_{3}\ \rangle\rvert$
@@
$p(a_{1}=-a_{2}=.....=-a_{N-1}=-a_{N}|A_{1},A_{2},....,A_{N})$$\geq$ $\frac{1}{2^{N-1}}\lvert\langle\ (A_{1}-A_{2}-A_{3}.....-A_{N-1}-A_{N})\ \otimes B_{2^{N-1}}\ \rangle\rvert$ (49)
@@
$p\Big{(}a_{1}=(-1)^{l_{y}^{2}}a_{2}=.....=(-1)^{l_{y}^{N-1}}a_{N-1}=(-1)^{l_{y}^{N}}a_{N}\ |\ A_{1},A_{2},....,A_{N}\Big{)}\geq\frac{1}{2^{N-1}}\lvert\sum_{x=1}^{N}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle\rvert.$ (50)
@@
$\sum_{x=1}^{N}\lvert\sum_{y=1}^{2^{N-1}}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle\rvert\leq 2^{N-1}$ (52)
By noting identity $|m+n|\leq|m|+|n|$, with $m,n$ being two arbitrary functions, we can write the above equation as,
$\lvert\sum_{x=1}^{N}\sum_{y=1}^{2^{N-1}}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle\rvert\leq 2^{N-1}$ (53)
@@
$\max_{B_{y}}\Big{(}\lvert\sum_{x=1}^{N}\sum_{y=1}^{2^{N-1}}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle\rvert\Big{)}\leq 2^{N-1}$ (54)
@@ Appendix B Existing necessary condition for joint measurability in quantum theory implied by our inequality
In order to discuss the necessary and sufficient condition for the $N$-wise compatibility of quantum measurements, let us take binary outcome observable $A_{x}$ where $A_{x}=\{E_{x}^{a_{x}}|E_{x}^{a_{x}}>0;\sum_{a_{x}}E_{x}^{a_{x}}=\mathbb{I}\}$. As discussed, $A_{x}$s are jointly measurable if there exists global POVMs $E_{1,2....N}^{a_{1},a_{2}...a_{N}}$ such that,
@@
It is already established [2, 3, 4] that for $N$ binary POVMs of the form, $E_{x}^{a_{x}}=\frac{\mathbb{I}+\eta a_{x}\Gamma_{x}}{2}$, where $\Gamma_{x}$ are the generators of Clifford algebra satisfying Eq. (40), the necessary condition for $N$-wise compatibility is,
@@
On the other hand, we derived the necessary condition for compatibility for any arbitrary operational theory as
$\eta\leq\frac{2^{N-1}}{\max_{B_{y}}\Big{(}\lvert\sum_{x=1}^{N}\sum_{y=1}^{2^{N-1}}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle\rvert\Big{)}}$ (57)
@@
Now, let us re-write our derived necessary condition for compatibility for $N=3$ as,

### v3

@@ VIII Summary and Conclusions
We note here that a number of studies have been made toward establishing a similar connection between incompatibility and various forms of quantum correlations, viz. nonlocality bene18 ; hirsch18 steering uola15 ; uola14 , preparation and measurement noncontextuality tavakoli20 ; singh23 . In this work, we have moved a step forward by establishing the connection between measurement incompatibility and generalised noncontextuality in any no-signalling operational theory, and thereby quantifying the maximum allowed incompatibility of a theory determined by the generalised contextuality.
@@
Furthermore, we have discussed the geometrical implications of the $N$-wise compatibility in a GPT. It has recently been proved in spekkens21 that a GPT is simplex-embeddable if and only if the ontological model corresponding to the operational theory this GPT concerns is generalised noncontextual. We demonstrated that any set of Alice’s observables that is used to remotely prepare Bob’s states is compatible if and only if Bob’s local statistics can be described in terms of a simplex-embeddable GPT. In rossi23 ; selby24 , it is shown that any preparation whose corresponding statistics cannot be described in terms of a simplex-embeddable GPT can be made simplex-embeddable by allowing dephasing noise to the states. In contrast to Refs.rossi23 ; selby24 , where prepare-measure scenario was considered, we argued that in our bipartite Bell scenario, Alice’s unsharp measurements remotely prepares noisy states so that a simplex-embeddable GPT explanation exists (which is otherwise not simplex-embeddable). Obviously, the degree of unsharpness parameter needs to be bounded. In quantum theory, simplex-embeddability exists for $\eta\leq 1/\sqrt{N}$.
@@
In recent work selby23c ; selby23 , it is argued that measurement incompatibility is not a prerequisite to reveal generalised contextuality in the preparation and measurement scenario. Although it apparently seems, there is no contradiction with the results of our paper. We claimed that the compatibility of Alice’s measurements used to steer Bob’s states is necessary and sufficient to reveal the generalised contextuality of the steered preparations. On the other hand, the claim of selby23c ; selby23 refers only to the connection between the measurements used to obtain the statistics from a single system and the possibility of revealing the generalised contextuality. Therefore, the domain of applicability of the results obtained in selby23c ; selby23 and this paper is different.
@@
Finally, we may consider a more general definition of the incompatibility of a theory in terms of the compatibility structure that exists in such a theory. In fact, it would be interesting to find connections between non-classical correlations exhibited by an arbitrary operational theory and other layers of non-classicality layers16 ; layers21 exist for measurements apart from the $N$-wise incompatibility. This would enable one to explore the uniqueness of quantum theory compared to other no-signaling theories. We note that, while discarding other class of incompatible theories in terms of our definition, we have assumed the superiority of generalized contextuality of quantum theory. It would be more intriguing if one can find a physical principle that can discard such theories and super-select the quantum theory. Such an attempt gonda18 was made which is particularly designed for the global joint measurability of sharp measurements through the Speaker’s principle. Studies along those direction could be an exciting avenue for future research and thus calls for further study.
@@
Acknowledgements:- S.M. acknowledges David Schmid for his comment on an earlier version of the manuscript. A.K.P acknowledges the support from the research grant SERB/MTR/2021/000908, Government of India. This work was also partly supported by Institute for Information and Communications Technology Planning and Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-$2023-00222863$).
@@ Appendix A Derivation of necessary condition for $N$-wise compatibility
$\ \ \ \ \ \ \ \ \ \ \ =\frac{1}{2^{N-1}}\langle\ (A_{1}+A_{2}+.....+A_{N})\ \otimes B_{1}\ \rangle-\Delta(N),$ (47)
where $\Delta(N)$ is a function of $N$-wise probabilities. Further, by using Eq. (45)-(A), we can write,
$p(a_{1}=a_{2}=.....=a_{N}|A_{1},A_{2},....,A_{N})\geq\frac{1}{2^{N-1}}\lvert\langle\ (A_{1}+A_{2}+.....+A_{N})\ \otimes B_{1}\ \rangle-\Delta(N)\rvert$ (48)
@@
$p(a_{1}=-a_{2}=.....=a_{N}|A_{1},A_{2},....,A_{N})$$\geq$ $\frac{1}{2^{N-1}}\lvert\langle\ (A_{1}-A_{2}+A_{3}.....+A_{N})\ \otimes B_{2}\ \rangle-\Delta(N)\rvert$
$p(a_{1}=a_{2}=-a_{3}=.....=a_{N}|A_{1},A_{2},A_{3}....,A_{N})$$\geq$ $\frac{1}{2^{N-1}}\lvert\langle\ (A_{1}+A_{2}-A_{3}.....+A_{N})\ \otimes B_{3}\ \rangle-\Delta(N)\rvert$
@@
$p(a_{1}=-a_{2}=.....=-a_{N-1}=-a_{N}|A_{1},A_{2},....,A_{N})$$\geq$ $\frac{1}{2^{N-1}}\lvert\langle\ (A_{1}-A_{2}-A_{3}.....-A_{N-1}-A_{N})\ \otimes B_{2^{N-1}}\ \rangle-\Delta(N)\rvert$ (49)
@@
$p\Big{(}a_{1}=(-1)^{l_{y}^{2}}a_{2}=.....=(-1)^{l_{y}^{N-1}}a_{N-1}=(-1)^{l_{y}^{N}}a_{N}\ |\ A_{1},A_{2},....,A_{N}\Big{)}\geq\frac{1}{2^{N-1}}\lvert\sum_{x=1}^{N}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle-\Delta(N)\rvert.$ (50)
@@
$\sum_{y=1}^{2^{N-1}}\lvert\sum_{x=1}^{N}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle-\Delta(N)\rvert\leq 2^{N-1}$ (52)
@@
$\max_{B_{y}}\Big{(}\sum_{y=1}^{2^{N-1}}\lvert\sum_{x=1}^{N}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle-\Delta(N)\rvert\Big{)}\leq 2^{N-1}$ (53)
@@ Appendix B Existing necessary condition for joint measurability in quantum theory implied by our inequality
In order to discuss the necessary condition for the $N$-wise compatibility of quantum measurements, let us take binary outcome observable $A_{x}$ where $A_{x}=\{E_{x}^{a_{x}}|E_{x}^{a_{x}}>0;\sum_{a_{x}}E_{x}^{a_{x}}=\mathbb{I}\}$. As discussed, $A_{x}$s are jointly measurable if there exists global POVMs $E_{1,2....N}^{a_{1},a_{2}...a_{N}}$ such that,
@@
It is already established kunjwal14 ; kunjwal14g ; kunjwal20 that for $N$ binary POVMs of the form, $E_{x}^{a_{x}}=\frac{\mathbb{I}+\eta a_{x}\Gamma_{x}}{2}$, where $\Gamma_{x}$ are the generators of Clifford algebra satisfying Eq. (40), the necessary condition for $N$-wise compatibility is,
@@
On the other hand, we derived the necessary condition for compatibility of measurements of the above form ($\Delta(3)=0$) as,
$\eta\leq\frac{2^{N-1}}{\max_{B_{y}}\Big{(}\sum_{y=1}^{2^{N-1}}\lvert\sum_{x=1}^{N}(-1)^{l_{y}^{x}}\langle A_{x}\otimes B_{y}\rangle\rvert\Big{)}\Big{)}}$ (56)
@@
Now, let us re-write our derived necessary condition for compatibility for $N=3$ mutually-maximally incompatible observables ($\Delta(3)=0$) as,